sum of the ages of a son and his father is 35 and the product of their age is 150 . find their ages..

Answer by ewatrrr(24785)


Hi
sum of the ages of a son and his father is 35
Let x represent represents father's age and (35-x) the son's age
the product of their age is 150
x(35-x) = 150
35x - x^2 = 150
x^2 - 35x + 150 = 0
Check your numbers given